OpenStudy (anonymous):

find the other endpoint of the line segment with the given endpoint and mid point. Endpoint:(-2,-5), Midpoint:(2,-7)

OpenStudy (anonymous):

you can set up an equation to solve this. the formula for finding the midpoint of a line is \[\frac{ x _{2}+x _{1} }{ 2 }, \frac{ y _{2}+y _{1} }{ 2 }\]

OpenStudy (anonymous):

you just plug in the point you have to solve for the missing coordinate so -2+x/2 = 2 -5+y/2=-7

OpenStudy (anonymous):

solve for x and y and you have your missing coordinates got it?

OpenStudy (anonymous):

ok thank you, but how do you solve for x and and y? thats were i get stuck..

OpenStudy (anonymous):

this is just linear algebra. for -2+x/2=2 multiply both sides by 2 to get rid of the fraction so -4+x=4 then add 4 to both sides->x=8

OpenStudy (anonymous):

same process to find y

OpenStudy (anonymous):

ohh ok thank you for the help

OpenStudy (anonymous):

wait sorry messed up

OpenStudy (anonymous):

x is 6

OpenStudy (anonymous):

same process just messed up the calculations

OpenStudy (anonymous):

so -2+x=4 x=6

OpenStudy (anonymous):

did you see my mistake?

OpenStudy (anonymous):

i think you added 4 + 4 when u were adding on both sides and you got 8.

OpenStudy (anonymous):

I just forgot that -2+x was all over x

OpenStudy (anonymous):

sorry all over 2

OpenStudy (anonymous):

for x-coordinates \[\frac{ -2+x }{2 } = 2\] for y-coordinates \[\frac{ -5+y }{ 2 } = -7\]

OpenStudy (anonymous):

Solve for x and y..... which is the other endpoint of the line segment.